OpenStudy (anonymous):

What is the perimeter of the triangle below? 24.14 28.26 30 110

OpenStudy (anonymous):



OpenStudy (anonymous):

Okay, so you have a isosceles right triangle. If you recall that the pythagorean therom is A^2+B^2=C^2, with C being the hypotenuse, you should see that A and B are equal. This means that C will always be the square root of the sums of A and B. You know what C is, so use that to solve for A and B and add :)

OpenStudy (anonymous):

A^2+B^2=\[\sqrt{10}\] where A and B are equal

OpenStudy (anonymous):

bry i don't understand it at all my teacher's never said how to get it i have no clue how to get that of what you're saying i looked up the pythagorean therom and i couldn't undertstand any of it

OpenStudy (anonymous):

K, well the Pythagorean therom just states that the hypotenuse of a right triangle is the sum of the squares of the two sides. so A^2+B^2=C^2. In your problem A and B are the two sides you don't know, but you do know that they are equal. C will equal 10. Since you know that A and B are equal you can rewrite the formula as 2A^2=10^2 and then just solve for A and you will know the length of all of your sides.

OpenStudy (anonymous):

does that help at all?

OpenStudy (anonymous):